# Conjugacy properties of multivariate unified skew–elliptical distributions

### v2

@@ Abstract
The broad class of multivariate unified skew–normal (sun) distributions has been recently shown to possess important conjugacy properties. When used as priors for the vector of parameters in general probit, tobit, and multinomial probit models, these distributions yield posteriors that still belong to the sun family. Although such a core result has led to important advancements in Bayesian inference and computation, its applicability beyond likelihoods associated with fully–observed, discretized, or censored realizations from multivariate Gaussian models remains yet unexplored. This article covers such an important gap by proving that the wider family of multivariate unified skew–elliptical (sue) distributions, which extends suns to more general perturbations of elliptical densities, guarantees conjugacy for broader classes of models, beyond those relying on fully–observed, discretized or censored Gaussians. Such a result leverages the closure under linear combinations, conditioning and marginalization of sue to prove that this family is conjugate to the likelihood induced by general multivariate regression models for fully–observed, censored or dichotomized realizations from skew–elliptical distributions. This advancement enlarges the set of models that enable conjugate Bayesian inference to general formulations arising from elliptical and skew–elliptical families, including the multivariate Student’s $t$ and skew–$t$, among others.
@@
Bayesian statistics , multivariate unified skew–elliptical distribution , unified skew–normal distribution.
@@ 1 Introduction
Conjugacy is a central property within Bayesian statistics. A prior $\mathcal{p}({\bm{\beta}})$ for the vector of parameters ${\bm{\beta}}\in\mathcal{B}\subset~{}\mathbb{R}^{p}$ is conjugate to the likelihood $\mathcal{p}({\bf{y}}\mid{\bm{\beta}})$ of the observed data ${\bf{y}}\in\mathcal{Y}\subset\mathbb{R}^{n}$, if the induced posterior $\mathcal{p}({\bm{\beta}}\mid{\bf{y}})$ still belongs to the same class of distributions of the assumed prior. This important property implies that when $\mathcal{p}({\bm{\beta}})$ belongs to a known and tractable family, Bayesian inference under the induced posterior can also leverage the tractability of such a family, thereby circumventing the challenges that arise in Bayesian computation and inference under intractable posterior distributions. Despite the relevance of this property, identifying known and tractable conjugate priors for the likelihoods induced by commonly–used statistical models is often challenging. Remarkably, until recently, conjugacy in regression settings was mainly established for univariate or multivariate normal responses ${\bf{y}}$ with Gaussian priors for the coefficients ${\bm{\beta}}$, thus hindering potentials of conjugate Bayesian inference beyond this specific setting.
@@
To address the aforementioned gap, Durante [32] has recently shown that also general probit models admit conjugate priors, with these priors belonging to the known family of unified skew–normal (sun) distributions [5]. Such a class includes multivariate Gaussians as a special case and extends these symmetric distributions through the perturbation of the corresponding density via a factor that coincides with the cumulative distribution function of a multivariate normal, thereby inducing skewness. Crucially, suns have (i) a known normalizing constant and moment–generating function, (ii) admit a tractable stochastic representation, and (iii) preserve the closure under linear combinations, conditioning, and marginalization of the original multivariate Gaussians [5, 18, 6]. These properties facilitate Bayesian inference under the induced sun posterior and, consequently, have motivated rapid subsequent research to establish sun conjugacy for broader classes of models beyond classical probit representations. Relevant advancements along these lines include dynamic multivariate probit [38], multinomial probit [36], probit Gaussian processes [27], tobit models [4] and, more generally, any representation inducing likelihoods proportional to the kernel of a sun [4]. Such a latter result crucially includes also important skewed extensions of classical probit, multinomial probit, and tobit [e.g, 28, 56, 12, 23, 47, 21, 48, 44, 22, 40], along with earlier conjugacy results for the parameters of skew–normal distributions [11, 61, 26]; see also Fasano et al. [37] and Onorati and Liseo [55] for additional results in binary regression settings, and Durante et al. [33] for an extension of the Bernstein–von Mises theorem which clarifies the crucial role played by skewed extensions of multivariate Gaussians in Bayesian approximations and asymptotic theory.
@@
Although the above contributions substantially enlarge the class of commonly–implemented statistical models that admit conjugate priors, all these formulations are based on fully–observed, discretized or censored Gaussian or skew–normal representations. As discussed above, this class includes multivariate linear regression along with probit, multinomial probit, and tobit models, among others, thus covering a core subset of formulations that are generally employed in statistics. Nonetheless, in several applications, there is often interest in extensions of these representations which replace the Gaussian or skew–normal assumption for the error terms with alternative distributions. Popular examples in applications are generalizations of linear regression, probit, multinomial probit, and tobit models that rely on Student’s $t$ or skew–$t$ error terms to incorporate robustness [e.g., 64, 50, 2, 8, 51, 31, 41, 52, 49, 46, 60]. More generally, several important contributions [e.g., 58, 25, 45, 20, 65] have also focused on multivariate elliptical [35] and skew–elliptical [16, 24, 17, 34, 10, 1] distributions, which include multivariate Gaussians, skew–normals, Student’s $t$ and skew–$t$ as special cases, thereby providing a large class of practically–relevant models. However, despite the relevance of such a family, there is a lack of general, unified, and tractable solutions for Bayesian inference within these settings. This is arguably due to the fact that, to date, no general conjugacy results have been established for generic models arising from fully–observed, censored, or dichotomized realizations from elliptical and skew–elliptical distributions.
@@
Motivated by this discussion, we cover such a gap by proving that multivariate unified skew–elliptical (sue) distributions [5, 10] are conjugate priors to the above class of models. From a technical perspective, the derivation of this result is based on specifying a general joint sue distribution for the parameters ${\bm{\beta}}$ and the noise vector ${\bm{\varepsilon}}$ of the response ${\bf{y}}$, and then leveraging the closure under linear combination, marginalization and conditioning of the sue family to prove that both $\mathcal{p}({\bm{\beta}})$ and $\mathcal{p}({\bm{\beta}}\mid{\bf{y}})$ are sue, whenever $\mathcal{p}({\bf{y}}\mid{\bm{\beta}})$ is proportional to a suitable likelihood induced by a fully–observed, censored or dichotomized elliptical or skew–elliptical distribution. Such a focus on the joint distribution serves only as a technical strategy to identify, under a classical Bayesian setting, which sue priors are conjugate to specific likelihoods, thereby yielding sue posterior distributions $\mathcal{p}({\bm{\beta}}\mid{\bf{y}})\propto\mathcal{p}({\bm{\beta}})\mathcal{p}({\bf{y}}\mid{\bm{\beta}})$ via the standard application of the Bayes rule. These novel results are obtained within Section 3, leveraging both available and newly–derived sue properties outlined in Section 2. As discussed in Sections 3.13.3 (see Examples 16), these advancements include, as a special case, the conjugacy properties derived in Anceschi et al. [4] for suns, while extending these properties to other models of potential practical interest, such as, for example, generalizations of linear regression, probit and tobit models to Student’s $t$ or skew–$t$ error terms. For these formulations, we show that the corresponding conjugate priors are multivariate unified skew–$t$ (sut[10, 62]. Concluding remarks can be found in Section 4, where we also clarify that besides the practical consequences for some special cases of the general results in Section 3, the conjugacy properties we derive are of broader and independent interest in expanding the theoretical analysis of the sue family.
@@
2 General overview and properties of multivariate unified skew–elliptical (sue) distributions
@@
Sections 2.12.2 provide an overview of the sue family along with its special cases, whereas Section 2.3 comprises both available closure properties and newly–derived ones that are required to prove the novel conjugacy results within Section 3. To ease the presentation, in the following, we adopt a different notation between random variables and the associated realizations only when the distinction between the two is not clear from the context.
2.1 Multivariate unified skew–elliptical distributions
@@
Multivariate unified skew–elliptical (sue) distributions [e.g., 5, 10] arise from the perturbation of elliptical densities [e.g., 35], defined as
@@
where ${\bm{\xi}}\in\mathbb{R}^{m}$ denotes a location parameter, $\bm{\Omega}\in\mathbb{R}^{m\times m}$ corresponds to a symmetric positive–definite dispersion matrix, and $g^{(m)}(\cdot):\mathbb{R}^{+}\to\mathbb{R}^{+}$ characterizes the so–called density generator. Recalling Fang et al. [35], different choices of such a density generator $g^{(m)}$ yield a broad class of routinely–implemented elliptical densities, covering multivariate Gaussians, Student’s $t$, Cauchy, logistic and Laplace, among others. As such, for a generic vector $\bar{{\bf{z}}}$ from an elliptical distribution it is customary to adopt the general notation $\bar{{\bf{z}}}\sim\mbox{{EC}}_{m}({\bm{\xi}},\bm{\Omega},g^{(m)})$, with ${\bm{\xi}}$, $\bm{\Omega}$ and $g^{(m)}$ parameterizing such a distribution. Refer to Chapters 1–3 of Fang et al. [35] for an in–depth treatment of the multivariate elliptical family, including details on the definition and properties of the density generator $g^{(m)}$.
Due to its generality, the multivariate elliptical family has been the subject of substantial interest, including in developing broader classes of distributions that introduce skewness in the above representation. An important and comprehensive example in this direction is provided by the sue family [5, 10]. Leveraging a parameterization that agrees with the unified skew–normal (sun) sub–family introduced by Arellano-Valle and Azzalini [5], and with the general selection representation in Arellano-Valle et al. [7] (see also Equation 19 in Arellano-Valle and Genton [10] and Section 7.1.3 of Azzalini and Capitanio [18]) a random vector ${\bf{z}}\in\mathbb{R}^{m}$ has a multivariate unified skew–elliptical (sue) distribution, i.e., ${\bf{z}}\sim{\mbox{SUE}}_{m,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bar{\bm{\Gamma}},g^{(m+q)})$, if its density $\mathcal{p}({\bf{z}})$ can be expressed as
@@
where $f_{m}({\bf{z}}-{\bm{\xi}};\bm{\Omega},g^{(m)})$ corresponds to the previously–defined elliptical density — evaluated at ${\bf{z}}$ — with density generator $g^{(m)}$, location ${\bm{\xi}}\in\mathbb{R}^{m}$, and positive–definite dispersion matrix $\bm{\Omega}\in\mathbb{R}^{m\times m}$ with associated scales and correlations in $\bm{\omega}=\text{diag}(\bm{\Omega})^{1/2}\in\mathbb{R}^{m\times m}$ and $\bar{\bm{\Omega}}=\bm{\omega}^{-1}\bm{\Omega}\bm{\omega}^{-1}\in\mathbb{R}^{m\times m}$, respectively. In addition, $\bm{\tau}\in\mathbb{R}^{q}$ is a truncation parameter, ${\bm{\Delta}}\in\mathbb{R}^{m\times q}$ denotes a shape matrix, whereas $\bar{\bm{\Gamma}}\in\mathbb{R}^{q\times q}$ corresponds to a positive-definite dispersion matrix. Finally, $Q({\bf{z}})$ is a quadratic form defined as $Q({\bf{z}})=({\bf{z}}-{\bm{\xi}})^{\top}\bm{\Omega}^{-1}({\bf{z}}-{\bm{\xi}})\in\mathbb{R}^{+}$, whereas $g_{Q({\bf{z}})}^{(q)}(u)=g^{(m+q)}[u+Q({\bf{z}})]/g^{(m)}[Q({\bf{z}})]$ denotes the elliptical conditional density generator. In (1), the generic function $F_{q}(\,\cdot\,;\bm{\Sigma},g^{(q)})$ is responsible for inducing skewness and is defined as the $q$–dimensional centered elliptical cumulative distribution function having dispersion matrix $\bm{\Sigma}\in\mathbb{R}^{q\times q}$ and density generator $g^{(q)}$. As clarified in detail within Lemma 5, the above density includes as a special case the one of classical elliptical distributions, which can be obtained by setting $\bm{\tau}=\bm{0}$ and ${\bm{\Delta}}=\bm{0}$. Such a result highlights that $\bm{\tau}$ and ${\bm{\Delta}}$ play a crucial role in inducing skewness. Let us also emphasize that, in this article, the notation $\bar{\bm{\Gamma}}$ is used to denote the Pearson–correlation matrix defined as $\bar{\bm{\Gamma}}=\bm{\gamma}^{-1}\bm{\Gamma}\bm{\gamma}^{-1}$, where $\bm{\gamma}=\text{diag}(\bm{\Gamma})^{1/2}$.
@@
${\bf{z}}\stackrel{{\scriptstyle d}}{{=}}(\bar{{\bf{z}}}\mid\bar{{\bf{z}}}_{0}>\bm{0}),\quad\mbox{with}\quad\begin{bmatrix}\bar{{\bf{z}}}\\ \bar{{\bf{z}}}_{0}\end{bmatrix}\sim\mbox{{EC}}_{m+q}\left(\begin{bmatrix}{\bm{\xi}}\\ \bm{\tau}\end{bmatrix},\begin{bmatrix}\bm{\Omega}&\bm{\omega}{\bm{\Delta}}\\ {\bm{\Delta}}^{\top}\bm{\omega}&\bar{\bm{\Gamma}}\end{bmatrix},g^{(m+q)}\right),$ (2)
where $\bar{{\bf{z}}}_{0}>\bm{0}$ indicates the event “each component of $\bar{{\bf{z}}}_{0}$ is positive”. More specially, under the above representation, a direct application of the Bayes rule yields
$\mathcal{p}({\bf{z}})=f_{m}({\bf{z}}-{\bm{\xi}};\bm{\Omega},g^{(m)})\frac{\mathbb{P}(-\bar{{\bf{z}}}_{0}\leq\bm{0}\mid\bar{{\bf{z}}}={\bf{z}})}{\mathbb{P}(-\bar{{\bf{z}}}_{0}\leq\bm{0})},\qquad{{\bf{z}}}\in\mathbb{R}^{m}.$ (3)
Hence, from the closure under linear combination, marginalization and conditioning of elliptical distributions [35, 18], it follows that $-\bar{{\bf{z}}}_{0}\sim\mbox{{EC}}_{q}(-\bm{\tau},\bar{\bm{\Gamma}},g^{(q)})$ and $(-\bar{{\bf{z}}}_{0}\mid\bar{{\bf{z}}}={\bf{z}})\sim\mbox{{EC}}_{q}(-\bm{\tau}-{\bm{\Delta}}^{\top}\bm{\omega}\bm{\Omega}^{-1}({{\bf{z}}}-{\bm{\xi}}),\bar{\bm{\Gamma}}-{\bm{\Delta}}^{\top}\bm{\omega}\bm{\Omega}^{-1}\bm{\omega}{\bm{\Delta}},g_{Q({\bf{z}})}^{(q)})$, where $\bm{\omega}\bm{\Omega}^{-1}=\bar{\bm{\Omega}}^{-1}\bm{\omega}^{-1}$ and $\bm{\omega}\bm{\Omega}^{-1}\bm{\omega}=\bar{\bm{\Omega}}^{-1}$; see also Arellano-Valle et al. [7]. As a consequence, the two quantities $\mathbb{P}(-\bar{{\bf{z}}}_{0}\leq\bm{0})$ and $\mathbb{P}(-\bar{{\bf{z}}}_{0}\leq\bm{0}\mid\bar{{\bf{z}}}={\bf{z}})$ at the denominator and numerator of Equation (3) coincide with the cumulative distribution functions, evaluated at $\bm{\tau}$ and $\bm{\tau}+{\bm{\Delta}}^{\top}\bar{\bm{\Omega}}^{-1}\bar{\bm{\omega}}^{-1}({{\bf{z}}}-{\bm{\xi}})$, of the centered elliptical distributions $\mbox{{EC}}_{q}(\bm{0},\bar{\bm{\Gamma}},g^{(q)})$ and $\mbox{{EC}}_{q}(\bm{0},\bar{\bm{\Gamma}}-{\bm{\Delta}}^{\top}\bar{\bm{\Omega}}^{-1}{\bm{\Delta}},\smash{g_{Q({\bf{z}})}^{(q)}})$, respectively, i.e., $F_{q}(\bm{\tau};\bar{\bm{\Gamma}},g^{(q)})$ and $F_{q}[\bm{\tau}+{\bm{\Delta}}^{\top}\bar{\bm{\Omega}}^{-1}\bm{\omega}^{-1}({\bf{z}}-{\bm{\xi}});\bar{\bm{\Gamma}}-{\bm{\Delta}}^{\top}\bar{\bm{\Omega}}^{-1}{\bm{\Delta}},g_{Q({\bf{z}})}^{(q)}]$, as in (1).
Besides providing additional insights on the quantities defining the joint density $\mathcal{p}({\bf{z}})$ in (1), the selection representation in (2)–(3) is also useful to derive the cumulative distribution function $\mathcal{P}({\bf{z}})$ of the sue distribution having density as in (1). More specifically, since the sue random vector ${\bf{z}}\sim{\mbox{SUE}}_{m,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bar{\bm{\Gamma}},g^{(m+q)})$ admits the equivalent selection representation in Equation (2), it follows that
@@
Therefore, leveraging again the closure under linear combinations of elliptical distributions [35, 18], it is possible to derive the following closed–form expression for the cumulative distribution function $\mathcal{P}({\bf{z}})$
@@
Notice that the vector $\bar{\bf{z}}_{0}$ is often called the latent part of the distribution and $q$ is the latent dimension.

### v3

# Conjugacy properties of multivariate unified skew-elliptical distributions

@@ Abstract
The broad class of multivariate unified skew-normal (SUN) distributions has been recently shown to possess important conjugacy properties. When used as priors for the coefficients vector in probit, tobit, and multinomial probit models, these distributions yield posteriors that still belong to the SUN family. Although this result has led to advancements in Bayesian inference and computation, its applicability beyond likelihoods associated with fully-observed, discretized, or censored realizations from multivariate Gaussian models remains yet unexplored. This article covers such a gap by proving that the wider family of multivariate unified skew-elliptical (SUE) distributions, which extends SUNs to more general perturbations of elliptical densities, guarantees conjugacy for broader classes of models, beyond those relying on fully-observed, discretized or censored Gaussians. Such a result leverages the closure under linear combinations, conditioning and marginalization of SUE to prove that this family is conjugate to the likelihood induced by multivariate regression models for fully-observed, censored or dichotomized realizations from skew-elliptical distributions. This advancement enlarges the set of models that enable conjugate Bayesian inference to general formulations arising from elliptical and skew-elliptical families, including the multivariate Student’s $t$ and skew-$t$, among others.
@@
Bayesian statistics , multivariate unified skew-elliptical distribution , unified skew-normal distribution.
@@ 1 Introduction
Conjugacy is a central property within Bayesian statistics. A prior $\mathcal{p}({\bm{\beta}})$ for the vector of parameters ${\bm{\beta}}\in\mathcal{B}\subset\mathbb{R}^{p}$ is conjugate to the likelihood $\mathcal{p}({\bf{y}}\mid{\bm{\beta}})$ of the observed data ${\bf{y}}\in\mathcal{Y}\subset\mathbb{R}^{n}$, if the induced posterior $\mathcal{p}({\bm{\beta}}\mid{\bf{y}})$ still belongs to the same class of distributions of the assumed prior. This important property implies that when $\mathcal{p}({\bm{\beta}})$ belongs to a known and tractable family, Bayesian inference under the induced posterior can also leverage the tractability of such a family, thereby circumventing the challenges that arise in Bayesian computation and inference under intractable posterior distributions. Despite the relevance of this property, identifying known and tractable conjugate priors for the likelihoods induced by commonly-used statistical models is often challenging. Remarkably, until recently, conjugacy in regression settings was mainly established for univariate or multivariate normal responses ${\bf{y}}$ with Gaussian priors for the coefficients ${\bm{\beta}}$, thus hindering potentials of conjugate Bayesian inference beyond this specific setting.
@@
To address the aforementioned gap, Durante [32] has recently shown that also general probit models admit conjugate priors, with these priors belonging to the known family of unified skew-normal (SUN) distributions [5]. Such a class includes multivariate Gaussians as a special case, and extends these symmetric distributions through the perturbation of the corresponding density via a factor that coincides with the cumulative distribution function of a multivariate normal, thereby inducing skewness. Crucially, SUNs (i) have a known normalizing constant and moment-generating function, (ii) admit a tractable stochastic representation, and (iii) preserve the closure under linear combinations, conditioning and marginalization of the original multivariate Gaussians [5, 18, 6]. These properties facilitate Bayesian inference under the induced SUN posterior and, consequently, have motivated rapid subsequent research to establish SUN conjugacy for broader classes of models beyond classical probit representations. Relevant advancements along these lines include dynamic multivariate probit [38], multinomial probit [36], probit Gaussian processes [27], tobit models [4] and, more generally, any representation inducing likelihoods proportional to the kernel of a SUN [4]. Such a latter result crucially includes also important skewed extensions of classical probit, multinomial probit, and tobit [e.g, 28, 56, 12, 23, 47, 21, 48, 44, 22, 40], along with earlier conjugacy results for the parameters of skew-normal distributions [11, 61, 26]; see also Fasano et al. [37] and Onorati and Liseo [55] for additional results in binary regression settings, and Durante et al. [33] for an extension of the Bernstein–von Mises theorem which clarifies the crucial role played by skewed extensions of multivariate Gaussians in Bayesian approximations and asymptotic theory.
@@
Although the above contributions substantially enlarge the class of routinely-implemented statistical models that admit conjugate priors, all these formulations are based on fully-observed, discretized or censored Gaussian or skew-normal representations. As discussed above, this class includes multivariate linear regression along with probit, multinomial probit, and tobit models, among others, thus covering a core subset of formulations that are often employed within statistics. Nonetheless, in several applications, there is still interest in extensions of these representations which replace the Gaussian or skew-normal assumption for the error terms with alternative distributions. Popular examples in applications are generalizations of linear regression, probit, multinomial probit, and tobit models that rely on Student’s $t$ or skew-$t$ error terms to incorporate robustness [e.g., 64, 50, 2, 8, 51, 31, 41, 52, 49, 46, 60]. More generally, several important contributions [e.g., 58, 25, 45, 20, 65] have also focused on multivariate elliptical [35] and skew-elliptical [16, 24, 17, 34, 10, 1] distributions, which include multivariate Gaussians, skew-normals, Student’s $t$ and skew-$t$ as special cases, thereby providing a large class of practically-relevant models. However, despite the relevance of such a family, there is a lack of general, unified and tractable solutions for Bayesian inference within these settings. This is arguably due to the fact that, to date, no general conjugacy results have been established for generic models arising from fully-observed, censored, or dichotomized realizations from elliptical and skew-elliptical distributions.
@@
Motivated by the above discussion, we cover such a gap by proving that multivariate unified skew-elliptical (SUE) distributions [5, 10] are the conjugate priors to the aforementioned class of models. From a technical perspective, the derivation of this result is based on specifying a general joint SUE distribution for the parameters ${\bm{\beta}}$ and the noise vector ${\bm{\varepsilon}}$ of the response ${\bf{y}}$, and then leveraging the closure under linear combination, marginalization and conditioning of SUE to prove that both $\mathcal{p}({\bm{\beta}})$ and $\mathcal{p}({\bm{\beta}}\mid{\bf{y}})$ are SUE, whenever $\mathcal{p}({\bf{y}}\mid{\bm{\beta}})$ is proportional to a suitable likelihood induced by a fully-observed, censored or dichotomized elliptical or skew-elliptical distribution. This focus on the joint distribution serves only as a technical strategy to identify, under a classical Bayesian setting, which SUE priors are conjugate to specific likelihoods, thereby yielding SUE posterior distributions $\mathcal{p}({\bm{\beta}}\mid{\bf{y}})\propto\mathcal{p}({\bm{\beta}})\mathcal{p}({\bf{y}}\mid{\bm{\beta}})$ via the standard application of the Bayes rule. These novel results are obtained within Section 3, leveraging both available and newly-derived SUE properties outlined in Section 2. As discussed in Sections 3.13.3 (see Examples 16), these advancements include, as a special case, the conjugacy properties derived in Anceschi et al. [4] for SUNs, while extending these properties to other models of potential practical interest, such as, for example, generalizations of linear regression, probit and tobit models to Student’s $t$ or skew-$t$ error terms. For these formulations, we show that the corresponding conjugate priors are multivariate unified skew-$t$ (SUT) [10, 62]. Concluding remarks can be found in Section 4, where we also clarify that besides the practical consequences for some special cases of the general results in Section 3, the conjugacy properties we derive are of broader and independent interest in expanding the theoretical analysis of the SUE family.
@@
2 General overview and properties of multivariate unified skew-elliptical (SUE) distributions
@@
Sections 2.12.2 provide an overview of the SUE family along with its special cases, whereas Section 2.3 comprises both available closure properties and newly-derived ones that are required to prove the novel conjugacy results within Section 3. To ease the presentation, in the following, we adopt a different notation between random variables and the associated realizations only when the distinction between the two is not clear from the context.
2.1 Multivariate unified skew-elliptical distributions
@@
Multivariate unified skew-elliptical (SUE) distributions [e.g., 5, 10] arise from the perturbation of elliptical densities [e.g., 35], defined as
@@
where ${\bm{\xi}}\in\mathbb{R}^{m}$ denotes a location parameter, $\bm{\Omega}\in\mathbb{R}^{m\times m}$ corresponds to a symmetric positive-definite dispersion matrix, and $g^{(m)}(\cdot):\mathbb{R}^{+}\to\mathbb{R}^{+}$ characterizes the so-called density generator. Recalling Fang et al. [35], different choices of such a density generator $g^{(m)}$ yield a broad class of routinely-implemented elliptical densities, covering multivariate Gaussians, Student’s $t$, Cauchy, logistic and Laplace, among others. As such, for a generic vector $\bar{{\bf{z}}}$ from an elliptical distribution it is customary to adopt the general notation $\bar{{\bf{z}}}\sim\mbox{{EC}}_{m}({\bm{\xi}},\bm{\Omega},g^{(m)})$, with ${\bm{\xi}}$, $\bm{\Omega}$ and $g^{(m)}$ parameterizing such a distribution. Refer to Chapters 1–3 of Fang et al. [35] for an in-depth treatment of the multivariate elliptical family, including details on the definition and properties of the density generator $g^{(m)}$.
Due to its generality, the multivariate elliptical family has been subject of a substantial interest that has led to the development of broader classes of distributions introducing skewness in the above representation. An important and comprehensive example in this direction is provided by the SUE family [5, 10]. Leveraging a parameterization that agrees with the unified skew-normal (SUN) sub-family introduced by Arellano-Valle and Azzalini [5], and with the general selection representation in Arellano-Valle et al. [7] (see also equation 19 in Arellano-Valle and Genton [10] and Section 7.1.3 of Azzalini and Capitanio [18]) a random vector ${\bf{z}}\in\mathbb{R}^{m}$ has a multivariate unified skew-elliptical (SUE) distribution, i.e., ${\bf{z}}\sim{\mbox{SUE}}_{m,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bar{\bm{\Gamma}},g^{(m+q)})$, if its density $\mathcal{p}({\bf{z}})$ can be expressed as
@@
where $f_{m}({\bf{z}}-{\bm{\xi}};\bm{\Omega},g^{(m)})$ corresponds to the previously-defined elliptical density — evaluated at ${\bf{z}}$ — with density generator $g^{(m)}$, location ${\bm{\xi}}\in\mathbb{R}^{m}$, and positive-definite dispersion matrix $\bm{\Omega}\in\mathbb{R}^{m\times m}$ with associated scales and correlations in $\bm{\omega}=\text{diag}(\bm{\Omega})^{1/2}\in\mathbb{R}^{m\times m}$ and $\bar{\bm{\Omega}}=\bm{\omega}^{-1}\bm{\Omega}\bm{\omega}^{-1}\in\mathbb{R}^{m\times m}$, respectively. In addition, $\bm{\tau}\in\mathbb{R}^{q}$ is a truncation parameter, ${\bm{\Delta}}\in\mathbb{R}^{m\times q}$ denotes a shape matrix, whereas $\bar{\bm{\Gamma}}\in\mathbb{R}^{q\times q}$ corresponds to a positive-definite dispersion matrix. Finally, $Q({\bf{z}})$ is a quadratic form defined as $Q({\bf{z}})=({\bf{z}}-{\bm{\xi}})^{\top}\bm{\Omega}^{-1}({\bf{z}}-{\bm{\xi}})\in\mathbb{R}^{+}$, whereas $g_{Q({\bf{z}})}^{(q)}(u)=g^{(m+q)}[u+Q({\bf{z}})]/g^{(m)}[Q({\bf{z}})]$ denotes the elliptical conditional density generator. In (1), the quantity responsible for inducing skewness is the $q$-dimensional centered elliptical cumulative distribution function $F_{q}(\,\cdot\,;\bar{\bm{\Gamma}}-{\bm{\Delta}}^{\top}\bar{\bm{\Omega}}^{-1}{\bm{\Delta}},g_{Q({\bf{z}})}^{(q)})$ with density generator $g_{Q({\bf{z}})}^{(q)}$, and dispersion matrix defined in the second argument. As clarified in Lemma 5, the density in  (1) includes as a special case the one of classical elliptical distributions, which can be obtained by setting $\bm{\tau}=\bm{0}$ and ${\bm{\Delta}}=\bm{0}$. Such a result highlights that $\bm{\tau}$ and ${\bm{\Delta}}$ play a crucial role in inducing skewness. Let us also emphasize that, in this article, the notation $\bar{\bm{\Gamma}}$ is used to denote the Pearson-correlation matrix defined as $\bar{\bm{\Gamma}}=\bm{\gamma}^{-1}\bm{\Gamma}\bm{\gamma}^{-1}$, where $\bm{\gamma}=\text{diag}(\bm{\Gamma})^{1/2}$.
@@
${\bf{z}}\stackrel{{\scriptstyle d}}{{=}}(\bar{{\bf{z}}}\mid\bar{{\bf{z}}}_{0}>\bm{0}),\qquad\qquad\begin{bmatrix}\bar{{\bf{z}}}\\ \bar{{\bf{z}}}_{0}\end{bmatrix}\sim\mbox{{EC}}_{m+q}\left(\begin{bmatrix}{\bm{\xi}}\\ \bm{\tau}\end{bmatrix},\begin{bmatrix}\bm{\Omega}&\bm{\omega}{\bm{\Delta}}\\ {\bm{\Delta}}^{\top}\bm{\omega}&\bar{\bm{\Gamma}}\end{bmatrix},g^{(m+q)}\right),$ (2)
where $\bar{{\bf{z}}}_{0}>\bm{0}$ indicates the event “each component of $\bar{{\bf{z}}}_{0}$ is positive”. More specifically, under the above representation and leveraging the closure under marginalization of elliptical distributions, a direct application of the Bayes rule yields
$\mathcal{p}({\bf{z}})=\mathcal{p}(\bar{{\bf{z}}}={\bf{z}})\frac{\mathbb{P}(-\bar{{\bf{z}}}_{0}\leq\bm{0}\mid\bar{{\bf{z}}}={\bf{z}})}{\mathbb{P}(-\bar{{\bf{z}}}_{0}\leq\bm{0})}=f_{m}({\bf{z}}-{\bm{\xi}};\bm{\Omega},g^{(m)})\frac{\mathbb{P}(-\bar{{\bf{z}}}_{0}\leq\bm{0}\mid\bar{{\bf{z}}}={\bf{z}})}{\mathbb{P}(-\bar{{\bf{z}}}_{0}\leq\bm{0})},\qquad{{\bf{z}}}\in\mathbb{R}^{m}.$ (3)
Letting $\bm{\omega}\bm{\Omega}^{-1}=\bar{\bm{\Omega}}^{-1}\bm{\omega}^{-1}$ and $\bm{\omega}\bm{\Omega}^{-1}\bm{\omega}=\bar{\bm{\Omega}}^{-1}$, and recalling again the closure properties of elliptical distributions [35, 18], it follows that $-\bar{{\bf{z}}}_{0}\sim\mbox{{EC}}_{q}(-\bm{\tau},\bar{\bm{\Gamma}},g^{(q)})$ and $(-\bar{{\bf{z}}}_{0}\mid\bar{{\bf{z}}}={\bf{z}})\sim\mbox{{EC}}_{q}(-\bm{\tau}-{\bm{\Delta}}^{\top}\bm{\omega}\bm{\Omega}^{-1}({{\bf{z}}}-{\bm{\xi}}),\bar{\bm{\Gamma}}-{\bm{\Delta}}^{\top}\bm{\omega}\bm{\Omega}^{-1}\bm{\omega}{\bm{\Delta}},g_{Q({\bf{z}})}^{(q)})$; see also Arellano-Valle et al. [7]. As a consequence, the two probabilities at the denominator and numerator of (3) coincide with the cumulative distribution functions, evaluated at $\bm{\tau}$ and $\bm{\tau}+{\bm{\Delta}}^{\top}\bar{\bm{\Omega}}^{-1}\bm{\omega}^{-1}({{\bf{z}}}-{\bm{\xi}})$, of the centered elliptical distributions $\mbox{{EC}}_{q}(\bm{0},\bar{\bm{\Gamma}},g^{(q)})$ and $\mbox{{EC}}_{q}(\bm{0},\bar{\bm{\Gamma}}-{\bm{\Delta}}^{\top}\bar{\bm{\Omega}}^{-1}{\bm{\Delta}},{g_{Q({\bf{z}})}^{(q)}})$, respectively, thereby allowing to recover (1).
Besides providing additional insights on the quantities defining the joint density $\mathcal{p}({\bf{z}})$ in (1), the selection representation in (2)–(3) is also useful to derive the cumulative distribution function $\mathcal{P}({\bf{z}})$ of the SUE distribution having density as in (1). More specifically, since the SUE random vector ${\bf{z}}\sim{\mbox{SUE}}_{m,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bar{\bm{\Gamma}},g^{(m+q)})$ admits the equivalent selection representation in (2), it follows that
@@
Therefore, leveraging again the closure under linear combinations of elliptical distributions [35, 18], it is possible to derive the following closed-form expression for the cumulative distribution function
@@
Notice that the vector $\bar{\bf{z}}_{0}$ is often called the latent part of the distribution, whereas $q$ is the latent dimension.